Allen operates a small sign-making business. He finds that if he charges x dollars for each sign, he sells 49 − x signs per week. What is the smallest number of signs that he can sell to have an income of $598 in one week?
Answer by lwsshak3(11628)

Allen operates a small sign-making business. He finds that if he charges x dollars for each sign, he sells 49 − x signs per week. What is the smallest number of signs that he can sell to have an income of $598 in one week?
***
price*quantity=income
let x=smallest number of signs that he can sell
x(49-x)=598
49x-x^2=598
x^2-49x+598=0
(x-23)(x-26)=0
x=23
or
x=26
smallest number of signs that he can sell=23